### v3

# UniCUE: Unified Recognition and Generation Framework for Chinese Cued Speech Video-to-Speech Generation

Abstract.
Cued Speech (CS) enhances lipreading via hand coding, offering visual phonemic cues that support precise speech perception for the hearing-impaired. The task of CS Video-to-Speech generation (CSV2S) aims to convert CS videos into intelligible speech signals. Most existing research focuses on CS Recognition (CSR), which transcribes video content into text. Consequently, a common solution for CSV2S is to integrate CSR with a text-to-speech (TTS) system. However, this pipeline relies on text as an intermediate medium, which may lead to error propagation and temporal misalignment between speech and CS video dynamics. In contrast, directly generating audio speech from CS video (direct CSV2S) often suffer from the inherent multimodal complexity and the limited availability of CS data. To address these challenges, we propose UniCUE, the first unified framework for CSV2S that directly generates speech from CS videos without relying on intermediate text. The core innovation of UniCUE lies in integrating a understanding task (CSR) that provides fine-grained CS visual-semantic cues to to guide the speech generation. Specifically, UniCUE incorporates a pose-aware visual processor, a semantic alignment pool that enables precise visual–semantic mapping, and a VisioPhonetic adapter to bridge the understanding and generation tasks within a unified architecture. To support this framework, we construct UniCUE-HI, a large-scale Mandarin CS dataset containing 11,282 videos from 14 cuers, including both hearing-impaired and normal-hearing individuals. Extensive experiments conducted on this dataset demonstrate that UniCUE achieves state-of-the-art (SOTA) performance across multiple evaluation metrics.
Chinese Cued Speech, Unified Framework, Video-to-Speech Generation, Understanding and Generation, Cued Speech Dataset
Figure 1. Illustration of the rules of the Chinese CS system and the proposed framework (UniCUE). (a) The chart for Mandarin Chinese CS (figure from (Liu and Feng, 2019)), where five distinct hand positions are used to encode vowels, and eight finger shapes are employed to represent consonants in Mandarin Chinese. (b) Our framework enables the direct generation of synchronized natural speech from video.
1. Introduction
Cued Speech (CS) is an visual phonetic encoding system that utilizes specific hand shapes and positions to enhance lip reading, providing an accurate visual representation of all phonemes in spoken language (Cornett, 1967; Liu and Feng, 2019; Leybaert and LaSasso, 2010). CS maintains a high level of consistency with spoken language in terms of phonemes and speech patterns, enabling hearing-impaired individuals to better integrate into speech-dominant social and educational environments (Cornett, 1967; Leybaert and LaSasso, 2010; Leybaert et al., 2010). In Mandarin Chinese, CS employs 8 hand shapes and 5 positions to encode consonants and vowels (as illustrated in Figure 1(a)), addressing challenges such as the phonemes with similar lip shapes (Liu and Feng, 2019).
CS Video-to-Speech generation (CSV2S) task aims to convert CS videos of into comprehensible speech signals. However, directly constructing an end-to-end CSV2S model faces several challenges. Firstly, this task involves complex multimodal semantic correlations, requiring precise mapping from visual cues (lip movements and hand coding) to acoustic speech, while the limited scale of existing CS datasets further constrains model capacity. Secondly, fine-grained spatiotemporal modeling of visual information is essential to resolve the intrinsic asynchrony, i.e., the hand-preceding phenomenon, where hand cues precede corresponding lip movements (Liu et al., 2020a). To the best of our knowledge, the CSV2S task has not been explicitly studied in prior literature.
Existing research primarily focuses on CS Recognition (CSR) that converts CS videos into phoneme-level text (Liu et al., 2020a, 2024, 2018, 2019), neglecting the critical need for natural speech generation. This limitation significantly impairs real-time communication between hearing-impaired and normal-hearing individuals, especially in educational and social scenarios. For instance, in group conversations, normal-hearing participants must quickly comprehend and respond to questions posed by their hearing-impaired peers. Textual output from CSR systems is often insufficient for such natural and smooth interactions. Additionally, recent lipreading-based video-to-speech models such as LipVoicer (Yemini et al., 2024) rely solely on lip movements, failing to capture the complementary hand-coded information in CS that conveys critical phonemic distinctions. These shortcomings underscore the need for a more comprehensive approach. Motivated by this, we aim to develop the first Chinese CSV2S system that directly decodes CS videos into intelligible speech, as illustrated in Figure 1(b).
A straightforward solution, shown in Figure 2(a), is to combine a CSR model with a Text-to-Speech (TTS) system. However, this combined pipeline suffers from two key drawbacks. Firstly, the intermediate textual representation introduces error propagation, as misrecognitions in the CSR stage lead to incorrect speech output. Secondly, the textual intermediate discards fine-grained spatiotemporal cues in the CS video, resulting in synthesized speech that lacks temporal alignment with the visual input.
Figure 2. (a) The combined CSV2S architecture combines separately trained CSR and TTS models. (b) Our unified framework (UniCUE) that transfers understanding capabilities of CSR into speech generation training by integrating the visual processor of CSR into CSV2S.
To overcome these challenges, we draw inspiration from recent advances in multimodal learning, where semantic reasoning from vision-language models (VLMs) has shown strong promise in tasks like text-guided image synthesis with interleaved control (Mi et al., 2025; Chen et al., 2025a). We hypothesize that the multimodal visual understanding inherent in CSR can serve as a semantic bridge to support more accurate and controllable speech generation in CSV2S. As depicted in Figure 2(b), we introduce a unified framework that leverages a shared visual processor to bridge CSR (understanding task) and CSV2S (generation task). This processor serves as a two-way translator: during CSR, it extracts linguistic semantics from fine-grained lip-hand motion patterns; in CSV2S, it utilizes these semantics to guide speech generation. The core innovation of our framework lies in modeling a semantic compensation flow, where phoneme-level supervision from CSR reduces ambiguity in speech synthesis, enabling more faithful and coherent voice generation under complex multimodal conditions.
Building upon this semantic compensation paradigm, in this work, we propose UniCUE, the first unified framework that bridges CSR and CSV2S tasks through three specific components: Firstly, unlike prior CSR methods (Liu et al., 2024; Liu and Liu, 2023) that process lip and hand modalities independently and rely on raw video embeddings, UniCUE employs a pose-aware processor that fuses video and pose streams into a mixed representation. This enables fine-grained spatiotemporal modeling of the hand-preceding phenomenon and improves generalization to cuer-specific expressive styles. Secondly, to enhance the alignment between visual and linguistic semantics, we introduce a semantic alignment pool to map the video and pose latent spaces into a shared textual space using contrastive learning. This facilitates cross-modal correlation modeling and improves semantic consistency in the generated speech. Thirdly, to unify the understanding and generation tasks, we reuse the CSR visual encoder within our diffusion-based CSV2S decoder and introduce a VisioPhonetic Adapter (VPA) that transforms the visual representations into diffusion-compatible codes. This design enables the decoder to effectively incorporate fine-grained semantic information derived from multimodal visual inputs
To evaluate UniCUE on hearing-impaired individuals, we extend the MCCS dataset (Lei et al., 2024b) by adding data from 8 hearing-impaired and 2 normal-hearing cuers [^1] Cuer means the people who perform CS., forming the Unified-HI Corpus with 14 cuers. Experimental results on this dataset demonstrate that UniCUE not only produces accurate and intelligible speech, but also maintains temporal synchronization with the CS video.
The main contributions of this work can be summarized as:
  • We propose the first CSV2S framework by constructing a unified multimodal system that integrates CSR capabilities to enhance speech generation.
  • We propose a pose-aware visual processor and a semantic alignment pool to enhance fine-grained, semantically aligned visual representations, and introduce an VPA module to convert fine-grained semantic information into understandable coding for the speech synthesis model.
  • We construct a new Mandarin Chinese CS dataset comprising both hearing-impaired and normal-hearing cuers. Experimental results demonstrate that our UniCUE outperforms the state-of-the-art (SOTA) methods in terms of speech accuracy, consistency, and quality.
Figure 3. Overview of our unified framework (UniCUE). It achieves direct Chinese CSV2S generation with semantic consistency, temporal alignment, and characteristics coherence by aligning the fine-grained spatiotemporal visual representations of CSR with the diffusion-based speech generator. The framework consists of three core modules: (1) Pose-Aware Visual Processor: Integrates video and pose embeddings to perform fine-grained spatiotemporal modeling of lip and hand movements. (2) Semantic Alignment Pool: Enhances the semantic mapping between visual features and speech content through video-text and pose-text contrastive learning. (3) VisioPhonetic Adapter (VPA): Converts fine-grained visual representation of CSR into condition encodings compatible with the diffusion-based generator.
2. Related Work
2.1. Video-to-Speech Generation
V2S aims to synthesize natural speech aligned with silent talking videos, but is challenged by limited data. Uni-Dubbing (Lei et al., 2024a) addresses this via modality-aligned pre-training on multimodal data and fine-tuning with both multimodal and audio-only inputs. Similarly, Kefalas et al. (Kefalas et al., 2024) pre-train on large audio-only corpora before tuning on paired data. Some studies (Kim et al., 2023; Yemini et al., 2024; Gupta et al., 2024) incorporate transcripts to enhance generation. Kim et al. (Kim et al., 2023) use text-speech supervision to improve word-level representation via multi-task learning. Existing V2S methods primarily focus on lip reading. However, CS conveys phonemic information through both lip and hand movements. Ignoring hand cues results in incomplete visual representations and degraded speech synthesis quality, limiting the applicability of these methods to CS. Notably, no prior work has addressed the CSV2S task.
2.2. Cued Speech Recognition
CS augments lip reading with hand coding to support the hearing-impaired. The CSR task aims to transcribe CS videos into text by leveraging lips and hands as complementary modalities (Liu and Feng, 2019; Papadimitriou and Potamianos, 2021). Most CSR methods extract lip and hand features separately and fuse them for recognition (Liu et al., 2020a, 2024; Liu and Liu, 2023; Zhang et al., 2023). Due to the asynchronous nature of these modalities, effective fusion remains challenging. Liu et al. (Liu et al., 2020a) proposed re-synchronization to align hand with lip features, while transformer-based mutual learning (Liu and Liu, 2023; Liu et al., 2024) improves multimodal interaction. Zhang et al. (Zhang et al., 2023) addressed privacy concerns via federated learning. In contrast, we directly model lip and hand cues from whole frames, avoiding explicit fusion. A pose-aware visual processor is introduced further to enhance cross-modal representation and improve performance.
2.3. Unified Understanding and Generation
Recent advances in unifying understanding and generation tasks fall into two main paradigms. The first integrates visual-language understanding with external generative models (e.g., diffusion models) for multimodal generation (Wu et al., 2024a; Dong et al., 2024; Jin et al., 2024; Li et al., 2024; Sun et al., 2024; Ge et al., 2024, 2025). For example, (Jin et al., 2024; Li et al., 2024) utilize large language models (LLMs) for semantic understanding and diffusion models (Rombach et al., 2022; Podell et al., 2023) for high-fidelity image synthesis. The second paradigm trains LLM-based foundation models via next-token prediction for both vision understanding and generation (Yu et al., 2023; Sun et al., 2023; Zhou et al., 2024; Wu et al., 2024b; Fang et al., 2024; Wu et al., 2025; Chen et al., 2025b). Transfusion (Zhou et al., 2024), for instance, unifies image understanding and generation within a single transformer, enabling controllable text-to-image synthesis by preserving visual details. However, existing approaches mainly focus on visual-text settings, leaving visual-to-speech generation underexplored. In this work, we introduce the first unified framework that bridges visual understanding and speech generation.
3. Method
3.1. Overview of UniCUE
To achieve accurate CSV2S generation, the proposed method needs to simultaneously address two critical challenges: (1) semantic understanding of the linguistic correlations between visual cues and speech content, and (2) speech synthesis that preserves cuer-specific characteristics and temporal alignment. Inspired by the auxiliary benefits of unified understanding and generation for multi-modal controllable image synthesis (Wu et al., 2024a; Dong et al., 2024), we design a unified architecture that integrates CSR and CSV2S, enabling CSV2S with understanding capability improvement through shared visual feature representations. As illustrated in Figure 3, the framework operates via two pathways.
CSR: Fine-grained Visual Cues Understanding. As the recognition pathway, CSR models fine-grained spatiotemporal visual semantics to transcribe CS videos into linguistic sequences. Given a CS video $I_{v}$ and its corresponding pose maps $I_{p}$ (extracted via OpenPose (Cao et al., 2019)), we first utilize a pose-aware visual processor to extract multi-modal embeddings $Z_{mv}$, which capture lip and hand motion cues. And then $Z_{mv}$ is fed into a auto-regressive Transformer-based text decoder $D_{T}$, which models long-range dependencies and contextual interactions across the sequence to generate the predicted token sequence: $T_{p}=D_{T}(Z_{mv})$, where $T_{p}$ denotes the predicted token sequence.
Unlike prior approaches relying on Connectionist Temporal Classification (CTC) loss (Graves et al., 2006), which predict each token independently and thus limit the model’s ability to capture cross-token dependencies and coarticulatory effects, our method employs an auto-regressive decoder $D_{T}$ supervised by cross-entropy loss. This design allows $D_{T}$ to generate tokens conditioned on previously generated outputs and spatialtemporal visual cues, which is more suited to modeling the asynchronous and dynamic nature of CS.
To further enhance both token-level precision and sequence-level linguistic consistency, we employ a hybrid training objective: a masked language modeling loss $\mathcal{L}_{\text{CE}}^{\text{masked}}$ supervises selectively masked ground-truth tokens to enhance contextual understanding; a sequence-level cross-entropy loss $\mathcal{L}_{\text{CE}}^{\text{seq}}$enforces supervision over the full sequence to promote accurate transcription. The final training objective for CSR is:
(1) $\mathcal{L}_{R}=\mathcal{L}_{\text{CE}}^{\text{masked}}(T_{p},T_{g})+\mathcal{L}_{\text{CE}}^{\text{seq}}(T_{p},T_{g}),$
where $T_{g}$ denotes the ground-truth token sequence. This dual-loss strategy enhances token-level accuracy while preserving global sequence semantics, enabling the model to capture subtle visual-linguistic cues and temporal dynamics inherent in CS videos, thus improving recognition performance and supporting speech synthesis.
CSV2S: Cuer-specific Speech Synthesis. To directly synthesize intelligible and personalized speech from CS videos, we formulate speech generation as a conditional denoising process within a latent diffusion model (LDM) (Rombach et al., 2022). Since both lip shapes and hand cues in CS convey phonemic content, the speech generation is conditioned a refined visual embedding $Z_{mv}^{\prime}$, which is derived by transforming the CSR multimodal feature $Z_{mv}$ via a VisioPhonetic adapter (VPA). Specifically, a pretrained VAE encoder compresses ground-truth mel-spectrograms into latent codes $Z_{s}$, which are progressively corrupted with Gaussian noise $\epsilon$ over $t$ steps: $Z_{s}^{t}:=\alpha_{t}\cdot Z_{s}+(1-\alpha_{t})\cdot\epsilon,$ where $\alpha_{t}$ denotes the noise level at timestep $t$. The noisy latent $Z_{s}^{t}$ then denoised by the LDM conditioned on $Z_{mv}^{\prime}$. The generation objective is defined as:
(2) $\mathcal{L}_{G}:=\mathbb{E}_{Z_{s}^{t},\,Z_{mv},\,\epsilon,\,t}\left[\left\|\epsilon-\mathcal{M}(Z_{s}^{t},Z_{mv},t)\right\|^{2}_{2}\right],$
where $\mathcal{M}$ represents the denoising network. By learning this conditional distribution, our model generates temporally aligned speech that reflects the visual expressions of cuers.
UniCUE: Unified Understanding and Generation. The CSR pathway learns fine-grained multi-modal visual embeddings $Z_{mv}$ through detailed linguistic recognition. To bridge the architectural gap between the CSR and the diffusion-based speech generator, we introduce a VPA that transforms $Z_{mv}$ into a refined representation $Z_{mv}^{\prime}$. These embeddings are subsequently utilized as conditional inputs to the CSV2S pathway, enabling the speech synthesis model to leverage enriched visual understanding for improved generation accuracy. By sharing visual feature representations within this unified framework, our approach effectively reduces information loss and mitigates error propagation that often arises from intermediate text conversions. As a result, CSV2S is capable of generating cue-specific speech that faithfully preserves linguistic fidelity and temporal alignment, producing personalized and intelligible speech outputs tailored to individual cuers.
3.2. Pose-aware Visual Processor
Considering the strong spatiotemporal correlation between hand coding, lip movement, and their underlying semantic content, both CSV2S and CSR require accurate modeling of lip and hand motion patterns. This necessitates a visual encoder capable of capturing fine-grained and temporally coherent features. While video frames offer rich appearance information, they often suffer from redundancy and visual ambiguity. In contrast, pose maps provide a compact, structured, and noise-resilient representation of motion dynamics. To leverage the complementary strengths of both modalities, we design a pose-aware visual processor that constructs fused visual representations, as shown in Figure 3.
Specifically, the input to the processor consists of video frames $I_{v}$ and pose maps $I_{p}$, both formatted as tensors of shape $T\times 3\times H\times W$, where $T$ indicates the frame lengths, and $H\times W$ denotes the spatial resolution. The processor comprises two main components. First, a shared visual encoder $E_{V}$ extracts spatiotemporal features from both modalities via a sequential architecture: a 2D ResNet backbone extracts frame-wise spatial features, which are stacked along the temporal axis and passed through a 1D temporal convolution to model short-term motion patterns. The resulting sequence is then fed into a Transformer encoder to capture long-range temporal dependencies across frames. This process yields the video features $Z_{v}=E_{V}(I_{v})$ and pose features $Z_{p}=E_{V}(I_{p})$, where $Z_{v}\in\mathbb{R}^{L\times D}$, $Z_{p}\in\mathbb{R}^{L\times D}$ with $D$ denoting the embedding dimension and $L=T\times N$ being the total number of tokens, where $N$ is the number of spatial patches per frame. Second, the projection layer integrates the two feature streams. The video and pose features are concatenated along the channel dimension and passed through a multi-layer perceptron (MLP), consisting of two linear layers with ReLU activation and LayerNorm, to produce the final mixed visual representation:
(3) $Z_{mv}=\text{MLP}(\text{Concat}(Z_{v},Z_{p})).$
The fused representation $Z_{mv}$ serves as a unified visual embedding that drives both recognition and generation pathways. In the subsequent modules, this representation is semantically aligned with linguistic content and refined for diffusion-based speech synthesis.
3.3. Semantic Alignment Pool
To further enhance semantic consistency between visual representation and linguistic content, we introduce a semantic alignment mechanism that aligns video, pose, and textual modalities through contrastive learning. Specifically, a ViT-based text encoder encodes the ground-truth transcript tokens $T_{g}$ into text embeddings $Z_{t}$. The visual features $Z_{v}$ and pose features $Z_{p}$, extracted by the pose-aware visual processor, are projected into a shared latent space via learnable linear layers. The text embedding $Z_{t}$ is similarly projected. We adopt a contrastive loss across the batch, treating each video-text and pose-text pair from the same sample as a positive pair, and all others as negatives. The loss is denoted as:
(4) $\mathcal{L}_{v\leftrightarrow t}=1-\cos(Z_{v},Z_{t}),\quad\mathcal{L}_{p\leftrightarrow t}=1-\cos(Z_{p},Z_{t}),$
where $\cos(\cdot,\cdot)$ denotes the cosine similarity between normalized embeddings. The total semantic alignment loss is calculated as:
(5) $\mathcal{L}_{S}=\mathcal{L}_{v\leftrightarrow t}+\mathcal{L}_{p\leftrightarrow t}.$
By enforcing this high-level alignment, the model is encouraged to extract complementary and discriminative semantics from visual modalities.These aligned features not only enhance linguistic recognition in CSR, but also offer semantically grounded condition for accurate speech synthesis.
Figure 4. The details of the VisioPhonetic Adapter, which transforms semantic visual embeddings into phonetic-aware features to enable seamless conditioning for diffusion-based speech synthesis.
3.4. VisioPhonetic Adapter
While the CSR-derived embeddings capture rich visual-linguistic semantics, they remain mismatched in format and granularity for direct use in diffusion-based speech generation. To bridge this modality gap, we propose the VisioPhonetic Adapter (VPA), which transforms semantically aligned visual features into a phonetic-aware conditioning signal suitable for the LDM. As illustrated in Figure 4, this lightweight module employs a sequential architecture to progressively refine visual-semantic representations into a diffusion-compatible conditioning signal:
(6) $\mathbf{Z}_{mv}^{\prime}=\text{MLP}\Big{(}\text{CrossAttn}\big{(}\text{MLP}(\mathbf{Z}_{mv})\big{)}\Big{)},$
which includes two MLPs and a Q-Former-style (Li et al., 2023) cross-attention layer. We use $N_{q}$ learnable semantic queries $\mathbf{f}\in\mathbb{R}^{N_{q}\times D}$, which is initialized by computing the average latent representation from ground-truth mel-spectrograms encoded by the pretrained VAE. This provides a phonetic-aware initialization aligned with the diffusion model’s target space. These queries act as phonetic slots to extract and reorganize relevant patterns from $\mathbf{Z}_{mv}$. The cross-attention mechanism operates as: $\mathbf{q}=\mathbf{W}^{q}\mathbf{f},\mathbf{k}=\mathbf{W}^{k}\mathbf{Z_{mv}},\mathbf{v}=\mathbf{W}^{v}\mathbf{Z_{mv}},\mathbf{a}=\text{Softmax}\left(\frac{\mathbf{q}\mathbf{k}^{T}}{\sqrt{d}}\right)\mathbf{v},\mathbf{{Z_{mv}}^{\prime}}=\text{MLP}(\mathbf{Z_{mv}}+a).$ The adapted features $\mathbf{Z}_{mv}^{\prime}$ serve as the final interface between visual understanding and speech synthesis, ensuring that the generated audio is not only temporally coherent but also linguistically faithful to the video input.
Table 1. Comparison between our Chinese Mandarin CS dataset and existing CS dataset. H denotes the cuers with normal hearing, while HI indicates hearing-impaired cuer. Our newly proposed Unified-HI Corpus is the first large-scale Chinese CS dataset with both hearing-impaired and normal-hearing cuers.
  Dataset Cuers Sentences Character Word Resolution FPS
  French CS (Liu et al., 2018) 1-H 238 12872 - $720\times 576$ 50
British CS (Liu et al., 2019) 1-H 97 2741 - $720\times 1280$ 25
MCCS (Lei et al., 2024b) 4-H 4000 131608 42256 $720\times 1280$ 30
Unified-HI (Ours)
6-H and 8-HI
11282 350333 112664 $720\times 1280$ 30
Table 2. Comparison with SOTA methods on test data of normal-hearing cuers and hearing-impaired cuers. Bold and underlined results are the best and second-best results. $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are preferable.
  Method Normal-hearing cuers Hearing-impaired cuers
[…]
  GT - 7.274 7.314 2.79 - - - - -
CMML 0.663 4.135 9.241 1.24 0.11 0.924 2.141 10.132 1.03
EcoCued 0.657 4.327 9.146 1.28 0.12 0.917 2.165 10.079 1.07
CSR (Ours) 0.186 4.874 9.125 2.53 0.57 0.224 3.342 9.315 2.29
  Lip2Speech 0.803 4.215 9.367 1.03 0.05 0.989 2.424 10.816 0.02
LipVoicer 0.754 4.361 9.226 1.12 0.08 0.971 2.623 10.517 0.04
CSV2S (Ours) 0.374 6.245 7.962 2.27 0.42 0.422 5.938 8.347 2.04
UniCUE (Ours) 0.205 6.729 7.632 2.46 0.53 0.248 6.491 8.076 2.17
4. Experiment
4.1. Experimental Setting
Dataset. Existing CS datasets are limited to normal-hearing cuers and lack data from hearing-impaired individuals, hindering model generalization to the primary users of assistive communication systems. To bridge this gap, we construct a new dataset, the Unified-HI Corpus, which includes CS videos from 8 hearing-impaired and 6 normal-hearing cuers. This diverse composition significantly enriches variations in gesture styles, lip movements, and speech patterns. The expanded coverage introduces more realistic challenges and better reflects practical use cases, enabling models to capture cue-specific nuances essential for hearing-impaired users. A comparison with existing CS datasets is shown in Table 1, and further details on sentence coverage and phoneme distribution are included in Appendix Section 2.
Due to the noisy speech data from hearing-impaired cuers, we use CS data from 6 normal-hearing cuers for training. The data from normal-hearing cuers is split by sentence into training and test sets with a 95:5 ratio to ensure effective training and validation. Importantly, all CS data from the 8 hearing-impaired cuers are used in the test set, enabling a robust evaluation of model generalization to this group.
Architecture Details. The CSV2S pathway is entirely built upon the AudioLDM  (Liu et al., 2023), including its VAE encoder-decoder, latent diffusion model, and vocoder components. For CSR, the Transformer in visual process, tokenizer, text-ViT, and text decoder are initialized from MBart (Liu et al., 2020b). Detailed training and inference configurations are provided in Appendix Section 1.
Evaluation Metrics. We evaluate the synthesized speech from three perspectives: linguistic accuracy, temporal synchronization, and speech quality. Linguistic accuracy is quantified by the Word Error Rate (WER) between the recognized text and ground truth. Temporal synchronization is assessed using SyncNet (Chung and Zisserman, 2016), reporting LSE-D (temporal distance) and LSE-C (confidence score). Speech quality is evaluated via STOI (Taal et al., 2010) for intelligibility and DNSMOS (Reddy et al., 2021) for naturalness.
Comparison Methods. We evaluate our UniCUE against: (1) CSV2S (Ours): direct speech synthesis without CSR assistance; (2) CSR (Ours): including pose-aware visual processor, text encoder and decoder, and semantic alignment pool; (3) CSR methods: CMML (Liu and Liu, 2023) and EcoCued (Liu et al., 2024); (4) V2S methods: Lip2Speech (Choi et al., 2023) and LipVoicer (Yemini et al., 2024).
4.2. Comparison with SOTA Methods
Quantitative Comparison. We compare our framework against SOTA methods, as summarized in Table 2. Our CSR model, empowered by the pose-aware visual processor and semantic alignment pool, achieves significantly lower WERs (0.186 for normal-hearing and 0.224 for hearing-impaired cuers), surpassing previous CSR methods. Building on this strong semantic understanding, UniCUE outperforms V2S methods across LSE-D, LSE-C, DNSMOS, and STOI metrics, demonstrating superior linguistic accuracy, temporal alignment, and speech quality.
Qualitative Comparison. Mel-spectrogram visualizations (Figure 4 in Appendix) further highlight the advantages of our method, showcasing improved temporal synchronization and clearer acoustic structures compared to others.
Table 3. Ablation Studies of model components on test data of norma hearing cuers and hearing-impaired cuers. The notations X††, X, and X* indicate ablated versions of the architecture X, where the pose maps, semantic alignment pool, and VPA module are removed, respectively.
  Method Normal-hearing cuers Hearing-impaired cuers
[…]
  GT - 7.274 7.314 2.79 - - - - -
CSR 0.210 4.746 9.129 2.42 0.49 0.250 3.218 9.402 2.19
CSR 0.204 4.783 9.224 2.46 0.53 0.247 3.234 9.397 2.21
CSR 0.186 4.874 9.125 2.53 0.57 0.224 3.342 9.315 2.29
CSV2S 0.398 6.158 8.122 2.21 0.40 0.398 5.821 8.582 1.96
CSV2S 0.374 6.245 7.962 2.27 0.42 0.422 5.938 8.347 2.04
UniCUE 0.239 6.637 7.724 2.30 0.44 0.267 6.419 8.163 2.08
UniCUE 0.231 6.641 7.716 2.33 0.46 0.276 6.421 8.159 2.10
UniCUE* 0.226 6.613 7.731 2.37 0.48 0.271 6.410 8.167 2.12
UniCUE 0.205 6.729 7.632 2.46 0.53 0.248 6.491 8.076 2.17
Figure 5. User study results for accuracy, quality, and synchronization metrics on normal-hearing (a) and hearing-impaired (b) test samples.
4.3. Ablation Studies
To verify the contribution of each component, we conduct ablation studies on both normal-hearing and hearing-impaired test data. Results are summarized in Table 3.
Unified Training Paradigm. Compared to direct CSV2S, UniCUE reduces WER by 45% (0.205 vs. 0.374) on normal-hearing cuers and 41% (0.248 vs. 0.422) on hearing-impaired cuers. These results highlight the benefit of leveraging fine-grained visual semantics from CSR to enhance CSV2S, alleviating the challenge of modeling complex multimodal correlations.
Visual Processor Design. Models that rely solely on raw video features struggle to capture fine-grained motion due to redundant and noisy visual information, resulting in suboptimal performance. By incorporating pose cues, our visual processor effectively captures cuer-specific dynamics, leading to significantly improved accuracy and robustness across diverse cuers. Semantic Alignment Mechanism. Disabling the Semantic Alignment Pool (SAP) degrades visual-semantic consistency, resulting in higher WERs for both CSR and UniCUE. This underscores the importance of the alignment in enforcing spatiotemporal coherence between visual cues and phonemic representations for accurate semantic modeling. The effectiveness of SAP is further validated by the t-SNE visualizations (Figure 5 in Appendix).
VisioPhonetic Adapter. Removing the VPA results in noticeable degradation in temporal alignment, demonstrating its crucial role in bridging the representation gap between CSR and CSV2S. By adaptively selecting and refining fine-grained spatialtemporal visual cues through learnable queries, the VPA enables more accurate and temporally coherent speech synthesis.
Impact of Hand Cues. Removing hand cues leads to substantial performance degradation, particularly for hearing-impaired users who often exhibit limited oral articulation and atypical lip shapes (see Appendix Table 1). The results highlight the complementary role of hand gestures in enhancing visual phonemic representations for CS.
Computational Efficiency. UniCUE achieves faster training convergence and 40% lower inference time than the combined pipeline (see details in Appendix Section 5).
4.4. User Study
To comprehensively assess the perceptual quality of synthesized speech, we conduct a user study involving 20 randomly selected test samples per cuer. Twenty volunteers rate the generated speech on three perceptual dimensions using 5-point Likert scales: Accuracy (1: unintelligible, 5: perfectly intelligible), Quality (1: artificial, 5: human-like), and Synchronization (1: desynchronized, 5: perfectly aligned). As shown in Figure 5, UniCUE consistently achieves significantly higher scores across all metrics, demonstrating statistically meaningful improvements. These findings validate that our unified framework effectively bridges visual understanding and speech generation, delivering superior performance in human perception compared to both modular pipelines and task-specific baselines.
5. Conclusion
This work introduces UniCUE, the first unified framework for directly generating speech from CS videos. By integrating fine-grained visual understanding with diffusion-based speech synthesis, UniCUE produces intelligible speech with precise temporal alignment. Key components including the pose-aware visual processor, semantic alignment pool, and VisioPhonetic Adapter, enable effective knowledge transfer from CS recognition (CSR) to CS video-to-speech generation (CSV2S), enhancing both linguistic accuracy and temporal synchronization. Additionally, we introduce the UniCUE-HI corpus, a new CS dataset featuring both normal-hearing and hearing-impaired cuers. Extensive experiments on this dataset demonstrate that UniCUE state-of-the-art methods across multiple evaluation metrics.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. 62471420), GuangDong Basic and Applied Basic Research Foundation (2025A1515012296), and CCF-Tencent Rhino-Bird Open Research Fund.